1) Cochrane, Inc., is considering a new three-year expansion project that requires an initial fixed asset investment of $2,370,000. The fixed asset will be depreciated straight-line to zero over its three-year tax life, after which time it will be worthless. The project is estimated to generate $2,240,000 in annual sales, with costs of $1,230,000.

Required:

If the tax rate is 35 percent, what is the OCF for this project? (Do not round intermediate calculations. Enter your answer in dollars, not millions of dollars (e.g., 1,234,567).)

  OCF $   

2) Cochrane, Inc., is considering a new three-year expansion project that requires an initial fixed asset investment of $2,250,000. The fixed asset will be depreciated straight-line to zero over its three-year tax life, after which time it will be worthless. The project is estimated to generate $2,200,000 in annual sales, with costs of $1,190,000. Assume the tax rate is 40 percent and the required return on the project is 11 percent.

Required:

What is the project
Answer #1

OCF = (Sales - Cost)(1 - tax rate) + (tax rate x Depreciation)

OCF = (2240000 - 1230000)*(1-0.35) + (0.35 X (2370000/3)

OCF = 656500 + 276500

OCF = $ 933000

2) OCF = (Sales - Cost)(1 - tax rate) + (tax rate x Depreciation)

OCF = (2200000 - 1190000)*(1-0.40) + (0.40 X (2250000/3)

OCF = 606000 + 300000

OCF = $ 906000

NPV = OCF x [1 - 1 / (1+r)^t / r] - initial investment

NPV = - $35994.47